OpenStudy (anonymous):

a telephone company charges $1.80 for a 12 minute long distance call, and $2.80 for a 20 minute call to the same number on the same day and time. find a linear equation that expresses the total charges c in terms of t.. @nincompoop @ayeedomo

OpenStudy (ayeedomo):

To find out how much the company charges each minute, divide $1.80 by 12: 1.80 / 12 = .15 So the telephone company charges 15 cents for each minute of the call ^ this is what I've calculated so far, BUT I'm a little stumped because the other call (2.80 for 20 minutes) amounts to 14 cents per minute...... Maybe I'm doing this completely wrong but I'll try to help XD

OpenStudy (anonymous):

i got c=t/8+2.80 because it is asking for the linear equation and i'm not sure if i did it right... thank you so much for your help..i really appreciate it

OpenStudy (anonymous):

i have one more last question for you if you don't mind me asking? @ayeedomo

OpenStudy (ayeedomo):

Sure :)

OpenStudy (anonymous):

a cellular phone company charges $10 for each phone plus $30 per month under a special sale plan. the total cost c (in dollars) of operating the phone is given by the linear model c(t)=10+30t, t is greater than or equal to 0 how many months would it take to have a total cost of $280

OpenStudy (ayeedomo):

For this question, what does t represent if c = total cost?

OpenStudy (ayeedomo):

Oh does t = the number of months?

OpenStudy (anonymous):

t is the in the equation that is given already....i was confused at that part too :(

OpenStudy (anonymous):

that question is a multiple choice question and the choices that are given are 6 months, 8 months, or 10 months!! i tired so hard on it but i'm not getting the right answer!

OpenStudy (ayeedomo):

Ok so: c(t)=10+30t c = total cost t = months(?) < If I'm wrong about that let me know, but the equation will be based on this equaling months So it says it costs 10 + 30(the number of months) Now if we fill in the info it should be: $280(t) =10+30(t)

OpenStudy (ayeedomo):

$280(t) =10+30(t) -10 -10 270(t) = 30(t) ^I'm pretty sure you're not supposed to calculate anything else for the left side, so now we're just calculating 270/30, which is 9... except 9 isn't an answer?!??!?!?

OpenStudy (anonymous):

i have an option that says none of the these.. do you think its that one because if we subtract 10 from both sides and than divide by 30 that will give us 9 right? but its not one of the options!

OpenStudy (ayeedomo):

There is a chance it could be none of these... But I don't want to make you get the question wrong so I'd try calculating it again, and if you still don't get any of those answers then it might be "none of these"

OpenStudy (ayeedomo):

Oh! I have an idea, hang on a second :)

OpenStudy (anonymous):

ok i'm here :)

OpenStudy (ayeedomo):

c(t)=10+30t So the answers available are 8, 6, and 10 months. We need to find the number of months that adds to a total of $280 $280 = 10+30t We could try plugging in the numbers: 10+30(8) = $250 10+30(6) = obviously lower than $250 lol 10+30(10) = $310 So I think it might be none of these after all

OpenStudy (anonymous):

i don't think it will work any way else :( i tried me many times solving it.. it is gotta be the none of these options

OpenStudy (anonymous):

Well, thank you so much for your big help...it was really nice working with you... i appreciated it..

OpenStudy (ayeedomo):

You're very welcome!! ^v^ Let me know if "none of these" turns out to be the right option :) Have a nice day!(or night lol) :D

OpenStudy (anonymous):

i'll let you know tomorrow for sure lol.... same to you as well :)

OpenStudy (anonymous):

so none of these was the correct answer :)

OpenStudy (ayeedomo):

YAY! XD